OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1 A. 69x + 10y = 51 B. -69x + 10y = -51 C. -69x + 10y = 51 D. 69x + 10y = -51 Could you help on this one please? @Hero a few moments ago

OpenStudy (anonymous):

it says that you need to express the answer as integers, an integer is a whole number, no decimals, and if you look at the answers, all they did was multiply everything by 10, now the standard form of a line is when you put both variables on one side and set that equal to the "y-intercept" which in the general form of a line is b\[y=mx+b\]so in standard form you are solving for that b

OpenStudy (anonymous):

thank you, is it D? @Embryo

OpenStudy (anonymous):

it should be A, the only thing you need to do is add the 6.9x to both sides so that it's on the same side as the y, then you multiply everything by 10 to get rid of the decimals, so everything should be positive

OpenStudy (anonymous):

thank you! i didnt even notice the negative 51. would you help with one more please? @Embryo

OpenStudy (anonymous):

sure, go ahead and post it up

OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 C. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (anonymous):

Alrighty, lets recall what slope intercept form is, it looks like this\[y=mx+b\]where m is the slope of the line, and b is the y-intercept, in order to go from the standard form to the slope intercept form, you need to solve for y, so add the 11x to both sides to get\[9y=11x-12\]then divide everything by 9 to get\[y=\frac{ 11 }{ 9 }x+-\frac{ 12 }{ 9 }=\frac{ 11 }{ 9 }x+-\frac{ 4 }{ 3 }\]now look at that and tell me what you get for b and m

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

that answer doesn't have the right negatives for the slope intercept form

OpenStudy (anonymous):

So its A?

OpenStudy (anonymous):

theyre the only two that make sense

OpenStudy (anonymous):

The slope is 11/9 and the y intercept is 4/3

OpenStudy (anonymous):

Okay, notice how i wrote the answer and how the slope-intercept form looks,\[y = (m)x + (b)\]\[y = (\frac{ 11 }{ 9 })x + (\frac{ -4 }{ 3 })\]now what is your "m" and what is your "b"

OpenStudy (anonymous):

m= 11/9 b= 4/3

OpenStudy (anonymous):

you're leaving out the negative for the 4/3, that negative is very important

OpenStudy (anonymous):

oh yeah! my bad!! b= -4/3

OpenStudy (anonymous):

there you go haha, never forget that the negative can't be ignored

OpenStudy (anonymous):

so b is my answer! lol thank you

OpenStudy (anonymous):

Yup, no problem